# Unlocking Unlabeled Data: Ensemble Learning with the Hui-Walter Paradigm for Performance Estimation in Online and Static Settings

### v1

@@ 2.2.2 Hui-Walter Online
Various algorithms work online, including stochastic gradient descent, k-means, and online support vector machines (Gomes et al., 2019). These algorithms differ in their update rules, but all aim to minimize the cumulative loss incurred over time as new data arrive. A critical aspect of online machine learning is the choice of the loss function, which determines the trade-off between the accuracy of the model and its ability to adapt to new data. However, these methods are still limited to using labeled data and assuming that ground truth is available. Online machine learning provides a powerful and flexible approach to real-time training machine learning models, but fails in some applied settings where the data is encrypted.
@@
A closed-form solution to the Hui-Walter estimates without a confidence interval appears in (Hui & Walter, 1980; Enùe et al., ). We reformulate our contingency table as follows:
@@
If this discriminant is zero or complex, then this online method cannot be calculated (Hui & Walter, 1980). One possible explanation for this phenomenon is the existence of Simpson’s paradox and the algebraic geometry that arises when dealing with three-way contingency tables (Slavkovi´c et al., 2009). The closed solutions to the Hui-Walter estimates are as follows:
@@
Using the formulas mentioned above, we can increase the streams’ counts and apply the estimates to the streaming data. This allows us to investigate the false positive and false negative rates, as well as the prior probabilities of streaming data at a specific time step. Interestingly, boosting provides an approach to this problem. It is a family of algorithms that convert multiple weak learners working in parallel into strong learners. This is where Hui-Walter online comes into play. With it, we can harness the power of these weak learners to assess the false-positive and false-negative rates, enhancing predictions by making them "prior aware" of the unlabeled data streams. Such an approach is not unique. In fact, several learning frameworks, such as the one mentioned by (Davis, 2020), advocate for better predictions by making the learner aware of previous information. However, a point to be taken into account is the sign of $F$ from equation 2. As pointed out by (Hui & Walter, 1980; Johnson et al., 2001), it can be either positive or negative, and sometimes, it does not present a solution at all. Hence, its value is determined by what yields plausible solutions, specifically within a probability simplex.
@@
1:procedure Online Hui-Walter($\mathcal{X},T$)
2:     $t\leftarrow 1$
3:     Initialize models $f_{1},f_{2}$ with $\mathcal{X}$
4:     Initialize three-way contingency table $\mathcal{T}$ and history $\mathcal{H}$
5:     while $t\leq T$ do
6:         Receive instance $x_{t}$
7:         Predict labels $y_{1}=f_{1}(x_{t}),y_{2}=f_{2}(x_{t})$
8:         Add $x_{t}$ to $\mathcal{H}$
9:         Get LPA profiles of $\mathcal{H}$
10:         Update cell counts of $\mathcal{T}$
11:         if t >> n then
12:              Calculate $\mathbf{X}=(\mathbf{\theta},\mathbf{\alpha},\mathbf{\beta}))$
13:         end if
14:         $t\leftarrow t+1$
15:     end while
16:end procedure
In Algorithm 2, $\mathcal{X}$ is the data set and $f_{1}$ and $f_{2}$ trained on subsets of the columns, $T$ is the number of instances to receive, $\mathcal{T}$ is a contingency table with as many dimensions as populations, $\mathcal{H}$ is the history of previously seen samples. The implementation of $\mathcal{H}$ could cache the previously seen samples in an LRU cache. The procedure calculates the latent profiles on the history $\mathcal{H}$ to assign the data to a population. After the $n$ instances have been predicted, the estimates for $\theta,\alpha,\beta$ start to be calculated and updated. The selection of $n$ should be reasonable such that there is enough stability in the values for $\theta,\alpha,\beta$ (lessening variation from one instance to the next) without sacrificing too many of the samples. This analogous to a "burn-in" concept commonly seen in MCMC (Hamra et al., 2013), but keep in mind that we are not discarding samples or assuming stationarity has been achieved.
3 Experimental Results
3.1 Data Sets
3.1.1 Wisconsin Breast Cancer
@@
The Wisconsin Breast Cancer Data Set has been widely used in machine learning and data mining to predict breast cancer diagnosis and prognosis. Researchers have used the data set to develop algorithms and models to classify tumors as benign or malignant and predict the probability of breast cancer recurrence or survival. These predictions can guide the treatment of patients with breast cancer and identify potential risk factors for the disease. However, it is essential to note that the data set has some limitations. For example, data is based on a specific population of patients and may not represent all cases of breast cancer. Furthermore, the data are from the 1990s and may not reflect more recent advances in breast cancer diagnosis and treatment. Despite these limitations, the Wisconsin breast cancer data set remains an essential resource for researchers studying breast cancer and working to improve the diagnosis and treatment of this disease. It has contributed significantly to our understanding of breast cancer and will continue to be a valuable resource for researchers in the field.
@@
In addition to its role in medicine, it has been used as an almost canonical data set for data science education and benchmarking of ML algorithms. It is one of the most downloaded data sets on the UCI Machine Learning Repository(Dua & Graff, 2017) [^1] https://archive.ics.uci.edu/ml/datasets/Breast+Cancer+Wisconsin+%28Diagnostic%29.
3.1.2 Adult
@@
This is another canonical data set from the UCI Machine Learning Repository(Dua & Graff, 2017)  [^2] https://archive.ics.uci.edu/ml/datasets/Adult. It contains census data from 1994 for the task of predicting whether an adult’s income exceeds \$50k/year. There are 32,561 observations described with a mix of categorical and continuous characteristics such as age, sex, education level, occupation, capital gains, and capital losses.
3.2 Hui-Walter Data Experiments
@@
3.2.1 Wisconsin Breast Cancer Data Set
Latent Profile Analysis
Figure 3: Five features from the Wisconsin Breast Cancer Data Set examined with Latent Profile Analysis
Figure 4: T-SNE dimension reduction of both latent populations.
We performed a Latent Profile Analysis (LPA) on the Wisconsin Breast Cancer data set to obtain two populations. Due to collinearities in the data set, a subset of features was selected to estimate the latent profiles. This subset was determined via hierarchical clustering on Spearman rank-order correlations and selecting one feature from each cluster, where clusters were defined as being separated by at least a distance of 1 per Ward’s linkage. LPA, assuming varying variances and varying covariances, was used to fit models assuming 1 through 10 latent classes. The approximate weight of evidence (AWE) criterion was used to identify the optimal number of classes, which was 2. AWE was used instead of BIC for model selection due to its ability to select more parsimonious models (Banfield & Raftery, 1993), especially relevant because we assumed complex parameterization (varying variances and varying covariances).
[…]
Figure 5: Bivariate plot of two Latent Profiles.
In Figure 5, we can see two distinct latent classes or subpopulations among the data. Figure 3 shows the columns of the data partitioned into one and then two latent classes. In Figure 4, we performed the t-SNE (Maaten & Hinton, 2008) dimension reduction of the columns of the Wisconsin Breast Cancer data set and labeled it according to the latent classes found with the analysis of the latent profile. We see that latent profile analysis and t-SNE separated the data into two latent populations.
Results
@@
Table 1: Model Predictions Over The Two Latent Populations
@@
Table 1 contains a three-way contingency table of categorizations on the Wisconsin Breast Cancer data set within each latent population based on the RF and SVC models. Both models predicted different proportions of malignant to benign tumors for the populations, suggesting models have some degree of independence.
With the models trained, we then proceeded to estimate parameters $\hat{\alpha}$, $\hat{\beta}$, and $\hat{\theta}$ without the ground truth labels by using the Hui-Walter framework with Gibbs Sampling. The sensitivity and specificity were each instantiated with a $Beta(1,1)$ distribution.
Table 2: Wisconsin Breast Cancer Data
@@
The estimated parameters using Hui-Walter with Gibbs Sampling are shown along with the actual values based on the ground truth in Table 2. For the SVC model, Hui-Walter estimated $\hat{\alpha}$ within 0.2 (0.101 compared to 0.132) and $\hat{\beta}$ within 0.01 (0.453 compared to 0.451). For the RF model, Hui-Walter estimates for $\hat{\alpha}$ was within 0.02 (0.044 compared to 0.031) and $\hat{\beta}$ was within 0.1 (0.310 compared to 0.211). The estimated priors $\hat{\theta}$ for Populations One and Two were close as well; Hui-Walter overestimated the prevalence in Population One (0.819 compared to 0.635) and underestimated for Population Two (0.035 compared to 0.122).
In Table 3, we have the confidence intervals for each parameter estimate. All the confidence intervals for the errors contain the parameter we found during the evaluation phase of the experiment. However, the true prevalence (prior or base rate) is outside the confidence interval for both populations ($\hat{\theta}_{1}$ and $\hat{\theta}_{2}$). Depending on context, these margins of error may be acceptable as a measure of classifier performance without requiring ground truth labeled data to compute.
Table 3: Estimated Parameters
@@
3.2.2 Adult
In lieu of latent profile analysis, we used the sex of the adult to partition the data set, resulting in two populations of $10,771$ females and $21,790$ males. We performed an $\frac{80}{20}$ split on the Adult data set to produce a training and test data set, sizes $26,048$ and $6,513$, respectively. We trained two classifiers, a linear regression (LR) and a random forest classifier (RF), on different columns of the data in the training data set.
Table 4: Model Predictions Over The Two Subpopulations
@@ 3.2.2 Adult
With the models trained, we then proceeded to estimate parameters $\hat{\alpha}$, $\hat{\beta}$, and $\hat{\theta}$ without the ground truth labels by using the Hui-Walter framework with Gibbs Sampling. The sensitivity and specificity were each instantiated with a $Beta(1,1)$ distribution.
Table 5: Adult Data Set
@@
Table 6: Estimated Parameters
@@
The Hui-Walter method obtained close estimates of some of the parameters. The $95\%$ confidence intervals for the rates of adults with over \$$\$50K$ income contained the ground truth rate for the male population (${\theta}_{2}$) and was 0.007 within the ground truth rate of the female population (${\theta}_{1}$). The confidence intervals for the error rate estimates ($\hat{\alpha}_{1}$, $\hat{\alpha}_{2}$, $\hat{\beta}_{1}$, $\hat{\beta}_{2}$) did not contain the actual values. However, $\hat{\alpha}_{1}$ and $\hat{\alpha}_{2}$ underestimated the actual by 7-10 times (0.008 compared to 0.058; 0.010 compared to 0.094) but $\hat{\beta}_{1}$ and $\hat{\beta}_{2}$ were closer, underestimating the actual by 2 times (0.392 compared to 0.554; 0.207 compared to 0.477).
3.2.3 Summary of Experimental Results

### v2

@@ 5 Hui-Walter Online
Various algorithms work online, including stochastic gradient descent [2], k-means [7], linear regression [1], and online support vector machines [17]. These algorithms differ in their update rules, but all aim to minimize the cumulative loss incurred over time as new data arrive. A critical aspect of online machine learning is the choice of the loss function, which determines the trade-off between the accuracy of the model and its ability to adapt to new data. However, these methods are still limited to using labeled data and assuming that ground truth is available. Online machine learning provides a powerful and flexible approach to real-time training machine learning models, but fails in some applied settings where the data is encrypted.
@@
A closed-form solution to the Hui-Walter estimates without a confidence interval appears in [19, 12]. We reformulate our contingency table in tabel 1 as follows:
@@
Table 1: Explanation of variables in terms of model agreement and disagreement indexed by population.
@@
If this discriminant is zero or complex, then this online method cannot be calculated [19]. One possible explanation for this phenomenon is the existence of Simpson’s paradox and the algebraic geometry that arises when dealing with three-way contingency tables [24]. The closed solutions to the Hui-Walter estimates are as follows:
@@
Using the formulas mentioned above, we can increase the streams’ counts and apply the estimates to the streaming data. This allows us to investigate the false-positive and false-negative rates, as well as the prior probabilities of streaming data at a specific time step. Interestingly, boosting provides an approach to this problem. It is a family of algorithms that convert multiple weak learners working in parallel into strong learners. This is where Hui-Walter online comes into play. With it, we can harness the power of these weak learners to assess the false-positive and false-negative rates, enhancing predictions by making them “prior aware” of the unlabeled data streams. This approach is not unique. In fact, several learning frameworks, such as the one mentioned in [8], advocate better predictions by making the learner aware of previous information. However, a point to be taken into account is the sign of $F$ from equation 2. As noted in [19, 22], it can be positive or negative, and sometimes it does not present a solution at all. Hence, its value is determined by what yields plausible solutions, specifically within a probability simplex.
@@
procedure Online Hui-Walter($\mathcal{X},T$)
     $t\leftarrow 1$
     Initialize models $f_{1},f_{2}$ with $\mathcal{X}$
     Initialize three-way contingency table $\mathcal{T}$ and history $\mathcal{H}$
     while $t\leq T$ do
         Receive instance $x_{t}$
         Predict labels $y_{1}=f_{1}(x_{t}),y_{2}=f_{2}(x_{t})$
         Add $x_{t}$ to $\mathcal{H}$
         Get LPA profiles of $\mathcal{H}$
         Update cell counts of $\mathcal{T}$
         if t >>much-greater-than$>$ >then
              Calculate $\mathbf{X}=(\mathbf{\theta},\mathbf{\alpha},\mathbf{\beta}))$
         end if
         $t\leftarrow t+1$
     end while
end procedure
Algorithm 2 Online algorithm to calculate model performance.
In Algorithm 2, $\mathcal{X}$ is the data set and $f_{1}$ and $f_{2}$ trained on subsets of columns of $\mathcal{X}$, $T$ is the number of instances to receive, $\mathcal{T}$ is a contingency table with as many dimensions as populations, $\mathcal{H}$ is the history of previously seen samples. After the $n$ instances have been predicted, the estimates for $\theta,\alpha,\beta$ start to be calculated and updated. The selection of $n$ should be reasonable such that there is enough stability in the values for $\theta,\alpha,\beta$ (lessening variation from one instance to the next) without sacrificing too many of the samples. This analogous to a “burn-in” concept commonly seen in MCMC [18], but keep in mind that we are not discarding samples or assuming stationarity has been achieved.
6 Data Sets
In this study, we used two canonical data sets, the Wisconsin Breast Cancer data set and the Adult data set, both commonly employed in machine learning research for benchmarking and classification tasks. We selected these data sets because of their well-documented attributes and applicability to evaluate the performance of classifiers. In the following, we provide an overview of each data set and the specific features used in our analysis.
6.1 Wisconsin Breast Cancer
@@
The Wisconsin Breast Cancer Data Set has been widely used in machine learning and data mining to predict breast cancer diagnosis and prognosis. Researchers have used the data set to develop algorithms and models to classify tumors as benign or malignant and predict the probability of breast cancer recurrence or survival. These predictions can guide the treatment of patients with breast cancer and identify potential risk factors for the disease.
@@
In addition to its role in medicine, it has been used as an almost canonical data set for data science education and benchmarking of ML algorithms. It is one of the most downloaded data sets on the UCI Machine Learning Repository[9] [^1] https://archive.ics.uci.edu/ml/datasets/Breast+Cancer+Wisconsin+%28Diagnostic%29.
Our adoption of a hierarchical clustering-based feature selection strategy [21] was a significant step in our efforts to reduce dimensionality and tackle multicollinearity in our data set. This approach, designed to eliminate redundant information and simplify the model, not only enhances interpretability and computational efficiency, but also plays a crucial role in preventing overfitting and improving the generalization performance of the model on unseen data [6].
We started by calculating the Spearman rank-order correlation matrix [33] for all characteristics. Spearman correlation effectively captures monotonic relationships between variables, accommodating linear and non-linear associations. This advantage makes it suitable for identifying linear and non-linear feature dependencies.
To prepare the data for clustering, we transformed the correlation matrix into a distance matrix [25] using the formula $D_{ij}=1-|\rho_{ij}|$, where $\rho_{ij}$ is the absolute Spearman correlation coefficient between features $i$ and $j$. This transformation ensures that highly correlated features—positively or negatively correlated—have smaller distances between them. As a result, features that share similar information are positioned closer together in the distance space.
Our use of Ward’s linkage method [34] in the hierarchical clustering process was instrumental in promoting the formation of compact and well-separated clusters. This method, by minimizing the total within-cluster variance, effectively identifies natural groupings in the data, where features within the same cluster are highly similar, and those in different clusters are dissimilar.
We formed clusters of the data by applying a distance threshold of $0.9$ to the hierarchical clustering dendrogram. This threshold grouped highly correlated features, effectively partitioning the set of characteristics into distinct clusters based on similarity. By setting this threshold, we controlled the granularity of the clustering, balancing retaining important features and reducing redundancy.
We selected one representative feature from each cluster [35] to include in our reduced feature set. The selection criterion was straightforward: we chose the first feature in each group. This approach simplifies the selection process, though alternative criteria could also be applied, such as selecting the feature with the highest variance or domain-specific importance. Following this procedure, we reduce the number of features from the original set to a smaller subset of representative features.
This approach selected five representative features: Mean Radius, Mean Texture, Mean Smoothness, Mean Compactness, and Texture Error (Texture SE). For each selected feature, we calculated the average intra-cluster and inter-cluster correlations to assess the effectiveness of our clustering method. The Mean Radius had an average intra-cluster correlation of $0.848\pm 0.186$, significantly higher than its average inter-cluster correlation of $0.338\pm 0.213$ ($p=2.237e-05$). Similarly, Mean Texture showed an average intra-cluster correlation of $0.909\pm 0.000$ versus an inter-cluster correlation of $0.254\pm 0.129$ ($p=3.448e-02)$. The other selected features also demonstrated higher intra-cluster correlations compared to inter-cluster correlations, with p-values indicating statistical significance: Mean Smoothness (intra: $0.541\pm 0.133$, inter: $0.308\pm 0.171$, $p=3.185e-03$), Mean Compactness (intra: $0.829\pm 0.053$, inter: $0.461\pm 0.183$, $p=2.330e-07$), and Texture SE (intra: $0.416\pm 0.027$, inter: $0.169\pm 0.119$, $p=2.217e-02$)  [^2] https://github.com/kslote1/hui-walter/blob/main/features.ipynb.
The cophenetic correlation coefficient of hierarchical clustering, which measures the degree of faithfulness the dendrogram maintains pariwise distances[13, 30], was $0.672$, indicating a moderate level of fidelity in representing pairwise feature dissimilarities. We then evaluated the performance of a Random Forest classifier using both the complete feature set and the reduced set of selected features. The model trained on all features achieved an accuracy of $95.6\%\pm 2.3\%$, while the model using the selected features achieved an accuracy of $92.3\%\pm 2.8\%$. The difference in model accuracy was not statistically significant ($p=0.065$), suggesting that the reduced feature set maintains comparable predictive power. These results demonstrate that our feature selection method effectively reduces dimensionality by eliminating redundant features while preserving the essential information necessary for accurate classification.
6.2 Adult
@@
This is another canonical data set from the UCI Machine Learning Repository[9]  [^3] https://archive.ics.uci.edu/ml/datasets/Adult. It contains census data from 1994 for the task of predicting whether an adult’s income exceeds \$50k/year. There are 32,561 observations described with a mix of categorical and continuous characteristics such as age, sex, education level, occupation, capital gains, and capital losses.
7 Hui-Walter Data Experiments
@@
7.1 Wisconsin Breast Cancer Data Set
Figure 3: Comparison of scaled features from the Wisconsin Breast Cancer Data Set distributions with and without latent classes. The top panel displays the distributions of five scaled features—Compactness Mean, Radius Mean, Smoothness Mean, Texture Mean, and Texture SE—segmented by latent profile classes. Individual data points are color-coded by class, with means, standard deviations, and 95% confidence intervals indicated. The bottom panel shows the same features aggregated without class separation, illustrating the overall distributions of the Latent Classes.
Figure 4: The two-dimensional t-SNE visualization of the training data from the Wisconsin breast cancer data set shows the viability of reducing the data space into Latent Classes. The high-dimensional feature space reduces to two dimensions using t-distributed Stochastic Neighbor Emulation (t-SNE)[26] with two components. Each point represents a single patient sample: benign cases are depicted as green squares and malignant cases are represented as red circles. This embedding reveals distinct clustering of benign and malignant samples, indicating that the selected features capture intrinsic differences between the two classes and suggest potential separability in the reduced dimensional space.
We performed a Latent Profile Analysis (LPA) on the Wisconsin Breast Cancer data set to obtain two populations. Due to collinearities in the data set, a subset of features was selected to estimate the latent profiles. This subset was determined via hierarchical clustering on Spearman rank-order correlations and selecting one feature from each cluster, where clusters were defined as being separated by at least a distance of 1 per Ward’s linkage. LPA, assuming varying variances and varying covariances, was used to fit models assuming 1 through 10 latent classes. The approximate weight of evidence (AWE) criterion was used to identify the optimal number of classes, which was 2. AWE was used instead of BIC for model selection due to its ability to select more parsimonious models [3], especially relevant because we assumed complex parameterization (varying variances and varying covariances).
Figure 5: A pairwise scatter plot matrix of selected morphological features from the Wisconsin Breast data set demonstrates the effectiveness of a two-class Latent Class Analysis (LCA) model. Plotted features are Radius Mean, Texture Mean, Smoothness Mean, Compactness Mean, and Texture SE, with data points colored by their assigned latent class. The distinct groupings in the plots indicate that the LCA effectively captures latent patterns in the data.
In Figure 5, to identify Latent Classes among the features, a pairwise scatter plot matrix of selected morphological features from the Wisconsin Breast shows the viability of identifying latent classes. The variables plotted are Radius Mean, Texture Mean, Smoothness Mean, Compactness Mean, and Texture SE. Each distribution represents a tumor feature, with colors indicating the assigned latent class (Latent Class 1 or Latent Class 2) from a two-class LCA model, a critical step in the process performed on binned categories of these variables. The diagonal panels display the density distributions of each feature, highlighting their univariate characteristics. The lower triangle panels present scatter plots between pairs of variables, illustrating bivariate relationships. The upper triangle panels show density plots of these feature pairs, providing additional insights into their joint distributions. Coloring by latent class reveals distinct groupings within the feature space, suggesting that the LCA effectively captures latent patterns in the data. Figure 3 shows the columns of the data partitioned into one and then two latent classes. In Figure 4, we performed the t-SNE [26] dimension reduction of the columns of the Wisconsin Breast Cancer data set and labeled it according to the latent classes found with the analysis of the latent profile. We see that latent profile analysis and t-SNE separated the data into two latent populations.
@@ 7.1 Wisconsin Breast Cancer Data Set
Table 1: Predictions from the model over the two latent populations.
Table 1 contains a three-way contingency table of categorizations on the Wisconsin Breast Cancer data set within each latent population based on the RF and SVC models. Both models predicted different proportions of malignant to benign tumors for the populations, suggesting that the models have some degree of independence.
With the models trained, we then proceeded to estimate parameters $\hat{\alpha}$, $\hat{\beta}$, and $\hat{\theta}$ without ground truth labels using the Hui-Walter framework with Gibbs sampling. Sensitivity and specificity were instantiated with a distribution of $Beta(1,1)$.
@@
Table 2: Top are the estimated parameters (TPR, TNR, prior) for both models. Below are the known values from the training environment in the Wisonsin Breast Cancer data Set.
The estimated parameters using Hui-Walter with Gibbs Sampling are shown along with the actual values based on the ground truth in Table 2. For the SVC model, Hui-Walter estimated $\hat{\alpha}$ within 0.03 (0.101 compared to 0.132) and $\hat{\beta}$ within 0.002 (0.453 compared to 0.451). For the RF model, the Hui-Walter estimates for $\hat{\alpha}$ were within 0.02 (0.044 compared to 0.031) and $\hat{\beta}$ was within 0.1 (0.310 compared to 0.211). The estimated priors $\hat{\theta}$ for Populations One and Two were close as well; Hui-Walter overestimated the prevalence in Population One (0.819 compared to 0.635) and underestimated for Population Two (0.035 compared to 0.122).
In Table 3, we have the confidence intervals for each parameter estimate. All confidence intervals for the errors contain the parameter we found during the evaluation phase of the experiment. However, the true prevalence (prior or base rate) is outside the confidence interval for both populations ($\hat{\theta}_{1}$ and $\hat{\theta}_{2}$). Depending on context, these margins of error may be acceptable as a measure of classifier performance without requiring ground truth labeled data to compute.
@@
Table 3: The estimated parameters and standard deviations from the Hui-Walter paradigm on static data for the Wisconsin Breast Cancer data set.
7.2 Adult
In lieu of latent profile analysis, we used the sex of the adult to partition the data set, resulting in two populations of $10,771$ females and $21,790$ males. We performed an $\nicefrac{{80}}{{20}}$ split on the Adult data set to produce a training and test data set, sizes $26,048$ and $6,513$, respectively. We trained two classifiers, a logistic regression (LR) and a random forest classifier (RF), on different columns of the data in the training data set.
@@ 7.2 Adult
With the models trained, we then proceeded to estimate parameters $\hat{\alpha}$, $\hat{\beta}$, and $\hat{\theta}$ without ground truth labels using the Hui-Walter framework with Gibbs sampling. Sensitivity and specificity were instantiated with a distribution of $Beta(1,1)$.
@@
Table 5: Top are the estimated parameters (TPR, TNR, prior) for both models. Below are the known values from the training environment in the Adult data set.
@@
Table 6: The estimated parameters and standard deviations from the Hui-Walter paradigm on static data for the Adult data set.
@@
The Hui-Walter method obtained close estimates of some of the parameters. The $95\%$ confidence intervals for the rates of adults with over \$$\$50\text{k}$ income contained the ground truth rate for the male population (${\theta}_{2}$) and was 0.007 within the ground truth rate of the female population (${\theta}_{1}$). The confidence intervals for the error rate estimates ($\hat{\alpha}_{1}$, $\hat{\alpha}_{2}$, $\hat{\beta}_{1}$, $\hat{\beta}_{2}$) did not contain the actual values. However, $\hat{\alpha}_{1}$ and $\hat{\alpha}_{2}$ underestimated the actual by 7-10 times (0.008 compared to 0.058; 0.010 compared to 0.094) but $\hat{\beta}_{1}$ and $\hat{\beta}_{2}$ were closer, underestimating the actual by 2 times (0.392 compared to 0.554; 0.207 compared to 0.477).
7.3 Summary of Experimental Results